OpenStudy (anonymous):

What is the simplified form of the expression 4the square root of v. + 4 the square root of v to the sixth power. − the square root of the quantity 25 times v. − the square root of the quantity 9 times v to the sixth power.?

OpenStudy (anonymous):

\[4\sqrt{ v}+4\sqrt{v^6}-\sqrt{25v}-\sqrt{9v^6}\] is this it?

OpenStudy (anonymous):

Yes, need the answer to it. @completeidiot

OpenStudy (anonymous):

@lgbasallote this person can help you

OpenStudy (lgbasallote):

who can help me?

OpenStudy (anonymous):

you can help elena- i wrote out the problem all nice and such

OpenStudy (lgbasallote):

if you want the "answer" then just go to wolframalpha, it's much faster. but if you want "help" then i'll help.

OpenStudy (anonymous):

anything please!